# Operator Learning for Continuous Spatial-Temporal Model with Gradient-Based and Derivative-Free Optimization Methods

### v2

@@ 3.3 Kuramoto–Sivashinsky Equation
To make the models better capture the long-term statistics, we retrain the models by jointly solving optimization problem Eq. (2.3) and  (2.4) with hybrid optimization training method described in Section 2.6. We take the pre-trained Model 3 (i.e., the model trained with the Resolution 3 in Table. 3.3) as an example and focus on the test performance with the same resolution, for which the long-term $D_{\textrm{KL}}$ in Table 3.3 shows a relatively large error. The forward map in Eq. (2.16) of EKI is constructed by the composition of three components: (i) long-term simulation (1000 time units) of the model with neural dynamical operator at the test resolution (i.e., $dx=22/256$ and $dt=2$), (ii) calculating the second-order spatial derivative $u_{xx}$ from the simulated system state $u$, and (iii) calculating the kurtosis of $u_{xx}$.
@@
Starting with the pre-trained Model 3 in Table 3.3, the hybrid optimization updates the neural dynamical operator by alternating the short-term trajectory matching via gradient-based optimization (i.e., Adam optimizer in this work) and the long-term statistics matching via derivative-free optimization (i.e., EKI in this work). More details about the hybrid optimization algorithm have been summarized in Algorithm 1. The training epochs of the gradient-based optimization is chosen as 3000 and the training epochs of EKI is chosen as 10, which means that all the trainable parameters in the neural dynamical operator will be updated by one EKI training epoch after every 300 epochs of the gradient-based optimization. The learning rate of the Adam optimizer is set as $10^{-4}$, considering that we are tuning a well-trained model at the beginning. In each EKI epoch, $N_{\textrm{it}}=20$ iterations will be performed with an ensemble size $J=100$.
@@
We present the error history during the EKI updating in Fig. 3.13. The short-term error is the mean squared error for short-term ($20$ time units) system state trajectory, while the long-term error is the mean squared error of the kurtosis of $u_{xx}$ from long-term ($1000$ time units) simulations. We can see that there is a trade-off between short-term state error and long-term statistics error in each EKI epoch, mainly because each EKI epoch only focuses on long-term statistics matching in the proposed hybrid optimization method. Although the short-term error tends to increase within each EKI epoch, the subsequent epochs of gradient-based optimization with short-term trajectory matching would keep tuning the neural dynamical operator such that a smaller short-term error is achieved. With the EKI loss training history, we selected the parameters trained after one iteration in $10^{\mathrm{th}}$ EKI epoch, highlighted by the black circle in the third column of Fig. 3.13.
@@ 4 Conclusion
A recent trend of data-driven modeling is to formulate the problem in its continuous form, which facilitates a more flexible use of data in general. The merits of existing spatially continuous models (e.g., operator learning) and temporally continuous models (e.g., neural ODE) have been demonstrated in many science and engineering applications. In this work, we present a data-driven modeling framework that learns a continuous spatial-temporal model based on the techniques of operator learning and neural ODE. More specifically, we focus on the learning of the dynamical operator and demonstrate that the learned model is resolution-invariance in both space and time. We also show that the learned model can provide stable long-term simulations, even if the training data only contains short-term time series of true system states. In addition, we propose a hybrid optimization scheme that leverages both gradient-based and derivative-free methods and efficiently combines the use of short-term time series and long-term statistics in training the model. The proposed framework is studied based on three classical examples governed by partial differential equations, including the viscous Burgers’ equation, the Navier–Stokes equation, and the Kuramoto–Sivashinsky equation. The results show that: (i) the trained model has resolution-invariance with respect to both spatial and temporal discretizations, and (ii) the hybrid optimization scheme ensures a good performance of the trained model in both matching short-term trajectories and capturing long-term system behaviors.
@@ Acknowledgments
The data that support the findings of this study are available from the corresponding author upon reasonable request. The codes and examples that support the findings of this study are available in the link: https://github.com/ChuanqiChenCC/Continous-Spatial-Temporal-Model.
@@ Appendix A Hybrid Optimization for Neural Dynamical Operator
1:$(u,t)\leftarrow$ (Training Data, Time Stamp)
2:$\tilde{\mathcal{G}}(\theta)\leftarrow$ Pre-trained Neural Operator by SGD with Short-term Data
3:$(T_{s},T_{l})\leftarrow(20,1000)$
4:$(N,k,\lambda,N_{\textrm{it}},J)\leftarrow(3000,300,1e^{-4},20,100)$
5:for  $i=1,2,...,N$  do $\triangleright$ SGD Training
6:     $t_{0}\sim\mathbb{U}(0,T-T_{s})$
7:     $u^{s}\leftarrow u[t_{0}:t_{0}+T_{s}]$
8:     $t^{s}\leftarrow t[t_{0}:t_{0}+T_{s}]$
9:     $\tilde{u}^{s}=\text{ODESolver}(\tilde{\mathcal{G}}(\theta),u^{s}[0],t_{0},t^{s})$
10:     $L^{s}=\frac{1}{T_{s}}||u^{s}-\tilde{u}^{s}||^{2}$
11:     $\theta\leftarrow\theta-\lambda\cdot\nabla L^{s}$triangleright Parameters updated by SGD
12:     if  $i\%k=0$  then $\triangleright$ EKI Training
13:         $t_{0}\sim\mathbb{U}(0,T-T_{l})$
14:         $u^{l}\leftarrow u[t_{0}:t_{0}+T_{l}]$
15:         $t^{l}\leftarrow t[t_{0}:t_{0}+T_{l}]$
16:         $y$ = stats($u^{l}$) $\triangleright$ Long-term Statistics
17:         $\mathbf{\Sigma}_{\eta}\leftarrow\text{diag}(0.01)$
18:         $\mathbf{\Theta}=\theta+\mathcal{N}(0,0.1^{2},(J,p))$
19:         for $n=1,2,...N_{\textrm{it}}$ do
20:              $\tilde{\mathbf{U}}=\text{ODESolver}(\tilde{\mathcal{G}}(\mathbf{\Theta}),u^{l}[0],t_{0},t^{l})$
21:              $\mathbf{G}=\text{stats}(\tilde{\mathbf{U}})$
22:              $\mathbf{Y}=y+\mathcal{N}(0,\bm{\Sigma}_{\eta},(J,))$
23:              $\mathbf{\Theta}\leftarrow\mathbf{\Theta}+[\mathbf{\Sigma}^{\theta g}(\mathbf{\Sigma}^{gg}+\mathbf{\Sigma}_{\eta})^{-1}(\mathbf{Y}-\mathbf{G})^{T}]^{T}$triangleright EKI updating formula
24:         end for
25:     end if
26:     $\theta\leftarrow\bar{\mathbf{\Theta}}$triangleright Parameters updated by EKI
27:end for

### v3

@@ 3.3.3 Hybrid Optimization for Short-Term State Prediction and Long-Term Statistics Matching
To make models better capture the long-term statistics, we take the pre-trained Model 3 (i.e., the model trained with the Resolution 3 in Table. 3.3) as an example and focus on the test performance with the same resolution, for which the long-term $D_{\textrm{KL}}$ in Table 3.3 shows a relatively large error. The forward map in Eq. (2.10) of EKI is constructed by the composition of three components: (i) long-term simulation (1000 time units) of the model with neural dynamical operator at the test resolution (i.e., $dx=22/256$ and $dt=2$), (ii) calculating the second-order spatial derivative $u_{xx}$ from the simulated system state $u$, and (iii) calculating the kurtosis of $u_{xx}$.
The main reason for starting from the pre-trained Model 3 is for efficiency consideration. More specifically, the long-term statistics engaged in this example is a scalar quantity, while the short-term loss aims to match a trajectory. Therefore, the amount of data is larger in the short-term loss than in the long-term one. Considering that the two losses are iteratively optimized in the hybrid algorithm, engaging the long-term loss at the beginning of the whole optimization process is not efficient and could have a negative impact on the convergence of the iterative optimization algorithm, as the information in the long-term statistics is far from enough to help guide the randomly initialized model to gradually mimic the true system. This example represents a large class of real-world applications, whose available long-term statistics are not as informative as the short-term trajectory data. In this work, we engage the long-term statistics via EKI as a modification of the pre-trained model with short-term loss only.
@@
Starting with the pre-trained Model 3 in Table 3.3, the hybrid optimization updates the neural dynamical operator by alternating the short-term trajectory matching via gradient-based optimization (i.e., Adam optimizer in this work) and the long-term statistics matching via derivative-free optimization (i.e., EKI in this work). More details about the hybrid optimization algorithm have been summarized in Algorithm 2. The number of training epochs $N_{\textrm{ep}}$ of the gradient-based optimization is chosen as 3000 and the number of training epochs of EKI is chosen as 10, which means that all the trainable parameters in the neural dynamical operator will be updated by one EKI training epoch after every 300 epochs of the gradient-based optimization. The learning rate of the Adam optimizer is set as $10^{-4}$, considering that we are tuning a well-trained model at the beginning. In each EKI epoch, $N_{\textrm{it}}=20$ iterations will be performed with an ensemble size $J=100$.
@@
We present the error history during the EKI updating in Fig. 3.13. The short-term error is the mean squared error for short-term ($20$ time units) system state trajectory, while the long-term error is the mean squared error of the kurtosis of $u_{xx}$ from long-term ($1000$ time units) simulations. We can see that there is a trade-off between short-term state error and long-term statistics error in each EKI epoch, mainly because each EKI epoch only focuses on long-term statistics matching in the proposed hybrid optimization method. Although the short-term error tends to increase within each EKI epoch, the subsequent epochs of gradient-based optimization with short-term trajectory matching would keep tuning the neural dynamical operator such that a smaller short-term error is achieved. With the EKI loss training history, we selected the parameters trained after one iteration in $10^{\mathrm{th}}$ EKI epoch, highlighted by the black circle in the third column of Fig. 3.13. It is worth noting that the proposed method sequentially optimizes the short-term loss $L_{s}$ and the long-term loss $L_{l}$ individually and then chooses a robust result, i.e., relatively small values of $L_{s}$ and $L_{l}$ compared to their range of values, respectively. In general, the approach is related to solving a two-objective optimization problem based on the Pareto front, which is estimated via iteratively solving the optimization problem that only involves $L_{s}$ or $L_{l}$.
@@ 4 Conclusion
A recent trend of data-driven modeling is to formulate the problem in its continuous form, which facilitates a more flexible use of data in general. The merits of existing spatially continuous models (e.g., neural operator) and temporally continuous models (e.g., neural ODE) have been demonstrated in many science and engineering applications. In this work, we present a data-driven modeling framework that learns a continuous spatial-temporal model based on the techniques of neural operator and neural ODE. More specifically, we focus on the learning of the dynamical operator and demonstrate that the learned model is resolution-invariance in both space and time. We also show that the learned model can provide stable long-term simulations, even if the training data only contains short-term time series of true system states. In addition, we propose a hybrid optimization scheme that leverages both gradient-based and derivative-free methods and efficiently combines the use of short-term time series and long-term statistics in training the model. The proposed framework is studied based on three classical examples governed by partial differential equations, including the viscous Burgers’ equation, the Navier–Stokes equation, and the Kuramoto–Sivashinsky equation. The results show that: (i) the trained model has resolution-invariance with respect to both spatial and temporal discretizations, and (ii) the hybrid optimization scheme ensures a good performance of the trained model in both matching short-term trajectories and capturing long-term system behaviors. Although this work mainly focuses on demonstrating the use of FNO to construct the neural dynamical operator and the iterative optimization of short-term and long-term losses, it is worth noting that other operator learning techniques (e.g., DeepONet) and optimization methods (e.g., $\epsilon-$constrained method for multiobjective optimization) can also be employed by the proposed framework.
@@ Acknowledgments
The data that support the findings of this study are available from the corresponding author upon reasonable request. The codes and examples that support the findings of this study are available in the link: https://github.com/AIMS-Madison/Neural-Dynamical-Operator.
@@ Appendix A Hybrid Optimization for Neural Dynamical Operator
1:Input:
2:  $\{\bm{u}(t_{n})\}_{n=0}^{N}\leftarrow$ Training Data
3:  $\{t_{n}\}_{n=0}^{N}\leftarrow$ Time Stamp
4:  $\tilde{\mathcal{G}}(\bm{\theta})\leftarrow$ Neural Operator
5:$(N_{s},N_{l})\leftarrow(20,1000)$
6:$(N_{\textrm{ep}},k,\lambda,N_{\textrm{it}},J)\leftarrow(3000,300,1e^{-3},20,100)$
7:for  $i=1,2,...,N_{\textrm{ep}}$  do $\triangleright$ SGD Training
8:     $N_{0}\sim\mathbb{U}(0,N-N_{s})$
9:     $\{\tilde{\bm{u}}(t_{n})\}_{n=N_{0}}^{N_{0}+N_{s}}=\text{ODESolver}(\tilde{\mathcal{G}}(\bm{\theta}),\bm{u}(t_{N_{0}}),\{t_{n}\}_{n=N_{0}}^{N_{0}+N_{s}})$
10:     $L_{s}=\Sigma_{n=N_{0}}^{N_{0}+N_{s}}||\bm{u}(t_{n})-\tilde{\bm{u}}(t_{n})||^{2}$
11:     $\bm{\theta}\leftarrow\bm{\theta}-\lambda\cdot\nabla_{\bm{\theta}}L_{s}$triangleright Parameters updated by SGD
12:     if  $i\bmod k=0$  then $\triangleright$ EKI Training
13:         $N_{0}\sim\mathbb{U}(0,N-N_{l})$
14:         $\bm{y}=\beta({\{\tilde{\bm{u}}(t_{n})\}_{n=N_{0}}^{N_{0}+N_{l}}})$triangleright Long-term Statistics
15:         $\bm{\Sigma}_{\bm{\eta}}=0.01\bm{\mathrm{I}}_{d_{\bm{y}}}$
16:         $\{\bm{\theta}^{(j)}\}_{j=1}^{J}=\bm{\theta}+\{\bm{\epsilon}^{(j)}\}_{j=1}^{J},\quad\bm{\epsilon}^{(j)}\sim\mathcal{N}(0,0.1^{2}\bm{\mathrm{I}}_{d_{\bm{\theta}}})$
17:         for $m=1,2,...N_{\textrm{it}}$ do
18:              {[$\{[\{\tilde{\bm{u}}(t_{n})\}_{n=0}^{N_{l}}]^{(j)}\}_{j=1}^{J}=\{\text{ODESolver}(\tilde{\mathcal{G}}(\bm{\theta}^{(j)}),\bm{u}(t_{N_{0}}),\{t_{n}\}_{n=N_{0}}^{N_{0}+N_{l}})\}_{j=1}^{J}$
19:              $\{\bm{g}^{(j)}\}_{j=1}^{J}=\{\beta([\{\tilde{\bm{u}}(t_{n})\}_{n=0}^{N_{l}}]^{(j)})\}_{j=1}^{J}$
20:              $\{\bm{y}^{(j)}\}_{j=1}^{J}=\bm{y}+\bm{\eta}^{(j)},\quad\bm{\eta}^{(j)}\sim\mathcal{N}(0,\bm{\Sigma}_{\eta})$
21:              $\{\bm{\theta}^{(j)}\}_{j=1}^{J}\leftarrow\{\bm{\theta}^{(j)}+\bm{\Sigma}^{\bm{\theta}\bm{g}}(\bm{\Sigma}^{\bm{gg}}+\bm{\Sigma}_{\bm{\eta}})^{-1}(\bm{y}^{(j)}-\bm{g}^{(j)})\}_{j=1}^{J}$triangleright EKI updating formula
22:         end for
23:     end if
24:     $\bm{\theta}=\frac{1}{J}\Sigma_{j=1}^{J}\bm{\theta}^{(j)}$triangleright Parameters updated by EKI
25:end for
26:return $\bm{\theta}$